# Kusatsu, Shiga 525-8577, Japan

### v1

@@
Gauged Courant sigma models Noriaki Ikeda ***E-mail: nikeda@se.ritsumei.ac.jp
Department of Mathematical Sciences, Ritsumeikan University
Kusatsu, Shiga 525-8577, Japan
@@ 1 Introduction
A Courant sigma model is a topological sigma model defined on a three-dimensional manifold [1, 2, 3], and it can be regarded as a generalization of Chern-Simons gauge theory. The target space of the model is a Courant algebroid $E$ [4] over a smooth manifold $M$. This model is also an example of AKSZ sigma models [5] constructed on mapping spaces between two graded manifolds and there are many mathematical and physical applications.
@@
In our previous work [6], we analyzed the gauging of the Poisson sigma model [7, 8] and the Dirac sigma model [9] by Lie groupoids (Lie algebroids) including Lie groups (Lie algebras). Target space geometry corresponds to geometry of Poisson manifolds with a group or groupoid actions. In that work, we formulated sigma model descriptions of Hamiltonian $G$-spaces and their generalizations, namely Hamiltonian Lie algebroid structures [10]. These models are two-dimensional AKSZ sigma models, that is, two-dimensional analogues of the Courant sigma model. They are referred to as the gauged Poisson sigma model (GPSM) and the gauged Dirac sigma model (GDSM).
The compatibility conditions between gauge symmetries and Poisson or Dirac structures are described by flatness conditions on curvature-type objects, such as the basic curvature of Lie algebroid connections. Moreover, when considering the GPSM and GDSM on manifolds with boundaries, boundary conditions are determined by Hamiltonian Lie algebroids [10] over Poisson manifolds [11] and Dirac structures [6]. Another approach to gauging the Poisson sigma model has been proposed in [12].
Since the Poisson sigma model is an example of an AKSZ sigma model, it is natural to consider the gauging of more general AKSZ sigma models. The Poisson sigma model is based on a QP-manifold of degree one. In this paper, we study the gauging of the Courant sigma model, which is an AKSZ sigma model based on a QP-manifold of degree two. For graded geometry including QP-manifolds, refer to [3, 13, 14, 15] and references therein.
Group-like objects of Courant algebroids are Lie $2$-groupoids. One should therefore consider actions of Lie $2$-groupoids, which serve as the appropriate group-like objects. See [16, 17, 18] for the integration of Courant algebroids to Lie $2$-groupoids. Another proposal for group-like objects associated with Courant algebroids is given by Lie rackoids [19]. In this paper, however, we restrict ourselves to infinitesimal data, namely Courant algebroids, since the Courant algebroid structure–or equivalently, a QP-manifold of degree two–is sufficient to construct an AKSZ action functional. The analysis of global group-like structures in the GCSM is left for future work.
@@
A Courant sigma model is constructed as an AKSZ sigma model, that is, it is based on a QP-manifold or, equivalently, a differential graded (dg) symplectic manifold. This construction relies on the fact that a Courant algebroid $E$ is in one-to-one correspondence with a QP-manifold of degree two, $T^{*}[2]E[1]$ [20]. By the AKSZ construction, a QP-manifold structure on the mapping space ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$ is induced from the QP-structure on $T^{*}[2]E[1]$. The resulting theory is equivalent to the Batalin-Vilkovisky formulation [21, 22] of the Courant sigma model. We employ the same construction for the gauged Courant sigma model in this paper.
To ensure the homological condition $Q^{2}=0$, where $Q$ is a degree-one vector field on the mapping space, certain geometric conditions corresponding to flatness of curvature-type objects are required. If such flatness conditions are not imposed, the theory should instead be regarded as an equivariant version of an AKSZ theory, in which $Q^{2}=0$ holds only up to groupoid actions. Such generalizations of AKSZ theories and equivariant versions of BV and AKSZ theories have been discussed in [23, 24, 25]. Gauged Courant sigma models provide concrete examples of this framework.
Moreover, in Section 4, and Apendix D, E and F, we consider deformations of gauged Courant sigma models by fluxes and boundary terms. By introducing various fluxes, consistency conditions are deformed. Conditions are regarded as generalizations of conditions of fluxes in string theory [26, 27].
If we consider boundary terms, the homological condition $Q^{2}=0$ imposes conditions of generalizations of momentum maps on boundary terms, including homotopy moment maps [28] and homotopy momentum sections [29] as special cases. As a result, we obtain generalizations of momentum maps to Courant algebroid settings.
@@ 2 Courant Sigma models
The action functional of a Courant sigma model [1, 3] is
@@
The space of fields in the Courant sigma model is the mapping space from $T[1]\Sigma$ to $T^{*}[2]E[1]$, ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$. Local coordinates on ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$ are called superfields. We denotes superfields by bold letters $(\bm{X}^{i}(\sigma,\theta),{\bm{Z}}_{i}(\sigma,\theta),{\bm{Y}}^{A}(\sigma,\theta))$ of total degree $(0,2,1)$, which correspond to local coordinates $(x^{i},z_{i},\eta^{A})$ on $T^{*}[2]E[1]$. Let $(\sigma^{\mu},\theta^{\mu})$ be local coordinates on $T[1]\Sigma$ of degree $(0,1)$. The theory has bidegree, degree in $T[1]\Sigma$ and degree in $T^{*}[2]E[1]$. Superfields are expanded by $\theta^{\mu}$ as
$\bm{X}^{i}(\sigma,\theta)$ $=X^{(0)i}(\sigma)+\theta^{\mu}X^{(1)i}_{\mu}(\sigma)+\frac{1}{2}\theta^{\mu}\theta^{\nu}X^{(2)i}_{\mu\nu}(\sigma)+\frac{1}{3!}\theta^{\mu}\theta^{\nu}\theta^{\lambda}X^{(3)i}_{\mu\nu\lambda}(\sigma),$ (6)
@@
$S_{C}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\rho_{i}^{A}(\bm{X}){\bm{Z}}_{i}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (9)
@@
$S_{C}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{Z}}_{i}{\bm{Q}}^{i}+\frac{1}{3!}H_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}\right],$ (10)
@@ 3.1 Standard Courant sigma models with Lie algebroid gauging
We introduce covariant local coordinates to calculate gauging by the Lie algebroid $A$. In order to make covariant QP-structures for the Lie algebroid action $A$, an ordinary connection $\nabla$ on $A$ is introduced. The corresponding connection $1$-form is $\omega=\omega^{a}_{bi}\mathrm{d}x^{i}\otimes e_{a}\otimes e_{b}$ such that $\nabla e_{a}=-\omega_{ai}^{b}\mathrm{d}x^{i}\otimes e_{b}$ for basis $e_{a}$ in $\Gamma(A)$For complete covariant expressions of the theory, we have to introduce an affine connection $\Gamma_{ij}^{k}$ on $TM$ for the standard Courant algebroid $TM\oplus T^{*}M$. However they are not introduced in this paper since the total QP-structure, the graded symplectic form and the homological function, do not depend on the affine connection. As the result, gauging conditions does not depend on an affine connection on $TM$, and only depend on a connection on $A$. . If a Lie algebroid $A$ is an action Lie algebroid $A=M\times\mathfrak{g}$, i.e., symmetries induced from a Lie algebra $\mathfrak{g}$, we do not need a connection since gauging does not depend on the connection.
@@
$\vartheta$ $=z_{i}\delta x^{i}+p_{i}\delta q^{i}+w_{a}\delta a^{a}$
$=z_{i}^{\nabla}\delta x^{i}-p_{i}\delta q^{i}+w_{a}Da^{a},$ (15)
@@
We can consider the structure is a equivariant QP-structure. A homological vector field $Q$ is extended to an equivariant vector field such that $Q^{\nabla}=Q+\iota_{\Omega}={\{{{\Theta^{\nabla}},{-}}\}}$ with a ’connection’ $\Omega$, where $Q^{\nabla}={\{{{\Theta^{\nabla}},{-}}\}}$. $(Q^{\nabla})^{2}=\iota_{{\mathcal{F}}}$ with a ’curvature’ $\mathcal{F}$}. Here $\mathcal{F}$} is the vector field induced from $(Q^{\nabla})^{2}$. The $Q^{\nabla}$-complex gives an equivariant cohomology up to Lie algebroid action. Equivariant versions of BV formalisms and AKSZ theories are analyzed in [24]. Physical theories with $Q^{2}\neq 0$ appear in double field theories [30, 31].
From the above data, an AKSZ action $S_{S}^{\nabla}$ is constructed. The kinetic term is induced from the Liouville form (15) as
$\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}\right],$ (30)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\bm{X}^{i}$. Other terms are induced from homological functions (21) and (22). As a covariant modification of Eq. (10), the covariantized deformations of the standard Courant sigma model is given by
$S_{S}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\frac{1}{3!}H_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}\right],$ (31)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{2}C_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right].$ (32)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}T_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right],$ (33)

### v2

# Kyoto 606-8502, Japan

@@
Abstract
@@ 1 Introduction
A Courant sigma model is a topological sigma model defined on a three-dimensional manifold Ikeda:2002wh ; Hofman:2002rv ; Roytenberg:2006qz , and it can be regarded as a generalization of Chern-Simons gauge theory. The target space of the model is a Courant algebroid $E$ LWX over a smooth manifold $M$. This model is also an example of AKSZ sigma models Alexandrov:1995kv constructed on mapping spaces between two graded manifolds and there are many mathematical and physical applications.
@@
In our previous work Ikeda:2023pdr , we analyzed the gauging of the Poisson sigma model Ikeda:1993fh ; Schaller:1994es and the Dirac sigma model Kotov:2004wz by Lie groupoids (Lie algebroids) including Lie groups (Lie algebras). Target space geometry corresponds to geometry of Poisson manifolds with a group or groupoid actions. In that work, we formulated sigma model descriptions of Hamiltonian $G$-spaces and their generalizations, namely Hamiltonian Lie algebroid structures Blohmann:2018 . These models are two-dimensional AKSZ sigma models, that is, two-dimensional analogues of the Courant sigma model. They are referred to as the gauged Poisson sigma model (GPSM) and the gauged Dirac sigma model (GDSM).
The compatibility conditions between gauge symmetries and Poisson or Dirac structures are described by flatness conditions on curvature-type objects, such as the basic curvature of Lie algebroid connections. Moreover, when considering the GPSM and GDSM on manifolds with boundaries, boundary conditions are determined by Hamiltonian Lie algebroids Blohmann:2018 over Poisson manifolds Blohmann:2023 and Dirac structures Ikeda:2023pdr . Another approach to gauging the Poisson sigma model has been proposed in Zucchini:2008cg .
Since the Poisson sigma model is an example of an AKSZ sigma model, it is natural to consider the gauging of more general AKSZ sigma models. The Poisson sigma model is based on a QP-manifold of degree one. In this paper, we study the gauging of the Courant sigma model, which is an AKSZ sigma model based on a QP-manifold of degree two. For graded geometry including QP-manifolds, refer to Roytenberg:2006qz ; Cattaneo:2010re ; Ikeda:2012pv ; Cueca:2025aej and references therein.
Group-like objects of Courant algebroids are Lie $2$-groupoids. One should therefore consider actions of Lie $2$-groupoids, which serve as the appropriate group-like objects. See LiBlandSevera ; MehtaTang1 ; MehtaTang2 for the integration of Courant algebroids to Lie $2$-groupoids. Another proposal for group-like objects associated with Courant algebroids is given by Lie rackoids LaurentGengouxWagemann . In this paper, however, we restrict ourselves to infinitesimal data, namely Courant algebroids, since the Courant algebroid structure–or equivalently, a QP-manifold of degree two–is sufficient to construct an AKSZ action functional. The analysis of global group-like structures in the GCSM is left for future work.
@@
A Courant sigma model is constructed as an AKSZ sigma model, that is, it is based on a QP-manifold or, equivalently, a differential graded (dg) symplectic manifold. This construction relies on the fact that a Courant algebroid $E$ is in one-to-one correspondence with a QP-manifold of degree two, $T^{*}[2]E[1]$ Roy01 . By the AKSZ construction, a QP-manifold structure on the mapping space ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$ is induced from the QP-structure on $T^{*}[2]E[1]$. The resulting theory is equivalent to the Batalin-Vilkovisky formulation BV1 ; BV2 of the Courant sigma model. We employ the same construction for the gauged Courant sigma model in this paper.
To ensure the homological condition $Q^{2}=0$, where $Q$ is a degree-one vector field on the mapping space, certain geometric conditions corresponding to flatness of curvature-type objects are required. If such flatness conditions are not imposed, the theory should instead be regarded as an equivariant version of an AKSZ theory, in which $Q^{2}=0$ holds only up to groupoid actions. Such generalizations of AKSZ theories and equivariant versions of BV and AKSZ theories have been discussed in Bonechi:2012kh ; Bonechi:2019dqk ; Grigoriev:2024ncm . Gauged Courant sigma models provide concrete examples of this framework.
Moreover, in Section 4, and Apendix D, E and F, we consider deformations of gauged Courant sigma models by fluxes and boundary terms. By introducing various fluxes, consistency conditions are deformed. Conditions are regarded as generalizations of conditions of fluxes in string theory Grana:2008yw ; Heller:2016abk .
If we consider boundary terms, the homological condition $Q^{2}=0$ imposes conditions of generalizations of momentum maps on boundary terms, including homotopy moment maps Callies:2013jbu and homotopy momentum sections Hirota:2021isx as special cases. As a result, we obtain generalizations of momentum maps to Courant algebroid settings.
@@ 2 Courant Sigma models
The action functional of a Courant sigma model Ikeda:2002wh ; Roytenberg:2006qz is
@@
The space of fields in the Courant sigma model is the mapping space from $T[1]\Sigma$ to $T^{*}[2]E[1]$, ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$. Local coordinates on ${\rm Map}(T[1]\Sigma,T^{*}[2]E[1])$ are called superfields. We denotes superfields by bold letters $(\boldsymbol{X}^{i}(\sigma,\theta),{\bm{Z}}_{i}(\sigma,\theta),{\bm{Y}}^{A}(\sigma,\theta))$ of total degree $(0,2,1)$, which correspond to local coordinates $(x^{i},z_{i},\eta^{A})$ on $T^{*}[2]E[1]$. Let $(\sigma^{\mu},\theta^{\mu})$ be local coordinates on $T[1]\Sigma$ of degree $(0,1)$. The theory has bidegree, degree in $T[1]\Sigma$ and degree in $T^{*}[2]E[1]$. Superfields are expanded by $\theta^{\mu}$ as
$\boldsymbol{X}^{i}(\sigma,\theta)$ $=X^{(0)i}(\sigma)+\theta^{\mu}X^{(1)i}_{\mu}(\sigma)+\frac{1}{2}\theta^{\mu}\theta^{\nu}X^{(2)i}_{\mu\nu}(\sigma)+\frac{1}{3!}\theta^{\mu}\theta^{\nu}\theta^{\lambda}X^{(3)i}_{\mu\nu\lambda}(\sigma),$ (6)
@@
$S_{C}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\rho^{i}_{A}(\boldsymbol{X}){\bm{Z}}_{i}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}\right],$ (9)
@@
$S_{C}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{Z}}_{i}{\bm{Q}}^{i}+\frac{1}{3!}H_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}\right],$ (10)
@@ 3.1 Standard Courant sigma models with Lie algebroid gauging
We introduce covariant local coordinates to calculate gauging by the Lie algebroid $A$. In order to make covariant QP-structures for the Lie algebroid action $A$, an ordinary connection $\nabla$ on $A$ is introduced. The corresponding connection $1$-form is $\omega=\omega^{a}_{bi}\mathrm{d}x^{i}\otimes e_{a}\otimes e_{b}$ such that $\nabla e_{a}=-\omega_{ai}^{b}\mathrm{d}x^{i}\otimes e_{b}$ for basis $e_{a}$ in $\Gamma(A)$ [^1] For complete covariant expressions of the theory, we have to introduce an affine connection $\Gamma_{ij}^{k}$ on $TM$ for the standard Courant algebroid $TM\oplus T^{*}M$. However they are not introduced in this paper since the total QP-structure, the graded symplectic form and the homological function, do not depend on the affine connection. As the result, gauging conditions does not depend on an affine connection on $TM$, and only depend on a connection on $A$. . If a Lie algebroid $A$ is an action Lie algebroid $A=M\times\mathfrak{g}$, i.e., symmetries induced from a Lie algebra $\mathfrak{g}$, we do not need a connection since gauging does not depend on the connection. One can refer to Appendix (A.2) for curvatures, $A$-connections, AA-torsions, basic curvatures, etc. in Lie algebroids.
@@
$\vartheta$ $=z_{i}\delta x^{i}-p_{i}\delta q^{i}-w_{a}\delta a^{a}$
$=z_{i}^{\nabla}\delta x^{i}-p_{i}\delta q^{i}-w_{a}Da^{a},$ (15)
@@
We can consider the structure is a equivariant QP-structure. A homological vector field $Q$ is extended to an equivariant vector field such that $Q^{\nabla}=Q+\iota_{\Omega}={\{{{\Theta^{\nabla}},{-}}\}}$ with a ’connection’ $\Omega$, where $Q^{\nabla}={\{{{\Theta^{\nabla}},{-}}\}}$. $(Q^{\nabla})^{2}={\mathcal{L}}_{{\mathcal{F}}}$ with a ’curvature’ $\mathcal{F}$}. Here $\mathcal{F}$} is the vector field induced from $(Q^{\nabla})^{2}$. The $Q^{\nabla}$-complex gives an equivariant cohomology up to Lie algebroid action. Equivariant versions of BV formalisms and AKSZ theories are analyzed in Bonechi:2019dqk . Physical theories with $Q^{2}\neq 0$ appear in double field theories Chatzistavrakidis:2018ztm ; Carow-Watamura:2018iau .
From the above data, an AKSZ action $S_{S}^{\nabla}$ is constructed. The kinetic term is induced from the Liouville $1$-form (15) as
$\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}-{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}-{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}-{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}-{\bm{W}}_{a}D{\bm{A}}^{a}\right],$ (30)
where $D{\bm{A}}^{a}=\mathrm{d}{\bm{A}}^{a}-\omega^{a}_{bi}{\bm{A}}^{b}\mathrm{d}\boldsymbol{X}^{i}$. Other terms are induced from homological functions (21) and (22). As a covariant modification of Eq. (10), the covariantized deformations of the standard Courant sigma model is given by
$S_{S}^{\nabla}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\frac{1}{3!}H_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}\right],$ (31)
@@
$S_{A}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{2}C_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right]$
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[{\bm{W}}_{a}D{\bm{A}}^{a}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right].$ (32)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}T_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right],$ (33)